# Affordance-based Robot Manipulation with Flow Matching

### v1

@@ 3.2.1 Flow Matching Model
     $\bm{x}_{t}=t\bm{x}_{1}+(1-t)\bm{x}_{0}$, interpolate trajectories
     $\mathbf{v}_{t}(\bm{x}|\bm{o}_{t})=f_{\bm{\theta}}(\bm{x}_{t},t|\bm{o}_{t})$, flow estimation
     $\nabla_{\bm{\theta}}\mathcal{L}_{\text{FM}}$, gradient step
@@
3.2.2 Flow Matching for Visuomotor Policy Learning
We extend flow matching to learn robot visuomotor policies. This requires two modifications in the formulation: i) changing the output $\bm{x}$ to represent robot actions; ii) modeling the flow estimation conditioned on input image observations. Fig. 4 illustrates our model structures.
In our case of robot manipulation, $\bm{x}_{1}$ in Equation (3) represents the demonstration robot trajectories. $\bm{x}_{0}$ is the random generated waypoints following a multivariate normal distribution $\bm{x}_{0}\sim\mathcal{N}(0,I)$. $\bm{x}$ here denotes the complete long-horizon robot trajectory for various tasks. We modify Equation (2) as:
$\mathbf{v}_{t}(\bm{x}|\bm{o}_{t})=f_{\bm{\theta}}(\bm{x}_{t},t|\bm{o}_{t})$
3.2.3 Implementation Details
The whole training process of the flow matching policy is illustrated in Algorithm 1 and Fig. 4. At time step $t$, the policy predicts flow vectors $\mathbf{v}_{t}$ for each waypoint conditioned on visual affordance observation data $\bm{o}_{t}$ with Feature-wise Linear Modulation (FiLM) (Perez et al., 2018) as well as the waypoints at the current timestep $\bm{x}_{t}$. The visual embeddings $\bm{o}_{t}$ are obtained through ResNet (He et al., 2016), and the flow model $f_{\bm{\theta}}$ is represented with U-Net. We will ablate various structures later.
For the inference procedure, random waypoints are sampled from the source distribution and then flowed into the target trajectory by estimating the flow from $t=0$ to $t=1$ over single or several steps:
$\bm{x}_{t+\Delta t}=x_{t}+\Delta tf(\bm{x}_{t},t|\bm{o}_{t}),\qquad\text{for}\ t\in[0,1]$ (4)
@@ 4 Experiments
We systematically evaluate prompt tuning method and flow matching policy against baseline studies. We also investigate how design choices would affect their performance.
We construct a real-world dataset with $10$ tasks across Activities of Daily Living to test the proposed method. Each task includes $1,000$ sets of RGB images, demonstrated robot trajectories, and labeled ground truth of affordances. The data has been manually collected by moving robot end-effectors with kinesthetic teaching. The novelty of our dataset could be summarized as: (i) It contains same scenarios with multiple objects, multi-task affordances and the demonstrated robot trajectories. (ii) All tasks are related to Activities of Daily Living that involve (simulated) human data. The dataset will be made publicly available soon. More details about tasks and prompts can be found in supplementary materials. We randomly split the dataset with $80$%-$20$% percentage of training and testing. The results reported here are obtained after $1,000$ epochs of training with 3-fold cross validation. The inputs and outputs are normalized for network training, which are then mapped back to the original scale for results reporting.
Our tasks include prompt primitives: ‘sweep the trash’, ‘pass the water to the human’, ‘hang the towel’, ‘put on the hat’, ‘cover the food’, ‘wipe the nose’, ‘wipe the forearm’, ‘feed the human’, ‘comb the hair’, and ‘brush the teeth’. Each task corresponds to one prompt primitive. We could also add a pretrained LLM layer (e. g., GPT) in the front for zero-shot text classification, allowing for linking other language instructions to one of the ten prompt primitives.
@@ 4.1 Affordance Evaluation with Prompt Tuning
4.1.1 Experiment Setup and dataset
We use a pretrained ViT-B-16 transformer as a vision backbone. The text layer adopts the classic setup, including Multiheaded Self-Attention, Feed-Forward Networks with LayerNorm and residual connections. As suggested by MAE (He et al., 2022), the decoder is only used for downstream tasks, and could be flexible and lightweight. Thus we use one single transformer decoder layer. For a fair comparison, all the baselines here use self-supervised pretrained MAE weights on ImageNet-21k dataset for the vision transformer model. In ablation studies, we will investigate different pretrained weights. The training parameters for the prompt tuning network include an image size of 224 × 224, AdamW with a learning rate of 1.5e-4 including Warmup with step-decay, and batch size of 256. The text encoding layers output 76 tokens, as the CLIP method (Radford et al., 2021).
@@ 4.1.1 Experiment Setup and dataset
4.1.2 Baseline Studies
@@
4.1.3 Main Results
Table 1 presents the results of prompt tuning on our testing dataset for affordance learning, comparing against baselines. We use two metrics to evaluate our results: (i) L2 error of affordance heatmap estimation (the fourth column), and (ii) L2 distance between the predicted and ground truth of heatmap centers (the fifth column). For the second metric, we label the heatmap affordance of our data with 2D Gaussian blobs centered on the pixel of demonstrated action during data collection. For the inference, we fit Gaussian Mixture Models to determine the heatmap centers. The heatmap error is averaged on each map, and the center error is averaged on per center point.
Four observations could be made from this result: (i) The deep structure outperforms the other baselines except for the full finetuning. (ii) Full finetuning slightly outperforms deep prompt tuning in terms of heatmap estimation error and heatmap center error. However, the distinction of heatmap center errors ($1.78$ pixels) remains subtle, given the full image size of $224\times 224$. This outcome is favorable as it indicates that most heatmap errors are caused by the tails of the Gaussian distribution, instead of the center area where the robot actions actually applied on. (iii) As expected, the shallow structure performs suboptimally to the deep structure. In the following ablation studies, we will investigate how the prompt depth affects the performance. (iv) The training dataset only includes a manikin. We observe that it generates well on our testing data with real humans.
What do prompts learn? We show a t-SNE (Van der Maaten and Hinton, 2008) visualization of the embeddings after the last vision transformer layer and before the decoder in Fig. 5. We can see that the points of the same color (e. g., tasks with same language prompts) are embedded together, implying that the representations recover the underlying manifold structure of discriminative task information.
@@ 4.2 Flow Matching Policy Evaluation
Figure 7: Training and testing loss of flow matching and diffusion policy throughout the training process. Flow matching exhibits greater stability on both training and evaluation than diffusion model.
@@ 4.2.1 Baseline Studies
We compare our flow matching policy against two other robot behavior cloning methods: (i) Diffusion Policy (Chi et al., 2023), and (ii) Transformer-based behavior cloning with Mean Square Error Loss, as customary in RVT (Goyal et al., 2023), RT-X (Padalkar et al., 2023).
Note that we are aware of other well-performed robot behavior cloning methods, including energy-based IBC (Florence et al., 2022), GAIL (Ho and Ermon, 2016), etc. Since extensive studies have been conducted between these methods and diffusion policy, we choose two representative baselines for evaluation. We consider CNN-based and Transformer-based structures for flow matching. The input of flow matching and baselines are the RGB image with visual affordance, and the output is a complete long-horizon trajectory in both 2D pixel space and 3D Cartesian space. All 10 tasks in our dataset are trained in a single policy with flow matching. For diffusion policy, we adopt the same setup as in the original paper (Chi et al., 2023), including $16$ inference denoising iterations for real-world robot testing.
@@
Baselines Diffusion Policy (16-step) 0.890 2.096 110.72
@@
Flow Matching (CNN, 1-step inference) 0.928 2.151 $8.53$
Ablations Flow Matching (CNN, 4-step inference) 0.890 2.094 34.99
Flow Matching (CNN, 8-step inference) 0.887 2.092 70.10
Flow Matching (CNN, 16-step inference) $0.886$ $2.091$ 113.68
@@ 4.2.2 Main Results
Table 2 presents the results of flow matching policy on our testing dataset for robot trajectory learning, comparing against baselines. We use two metrics for evaluation: (i) the error of 2D and 3D trajectory estimation, and (ii) the average inference time across various steps, performed using PyTorch with RTX 4090 GPU acceleration. The trajectory error is averaged on each point of the trajectory. We use 32 waypoints to represent the trajectory. In the later session, we will ablate this choice.
@@
Three observations could be made from this result: (i) Flow matching (CNN-based, 16 steps) outperforms other baselines in terms of 2D and 3D trajectory prediction accuracy. The 3D estimation performs decently with our method using only RGB images as input. But we also observe the 3D estimation could be highly attuned to camera angles. (ii) The Transformer behavior cloning achieves marginal precision. This is expected as it is hindered by the nature of multi-modal action distribution, causing the averaging out across non-convex spaces. (iii) Flow matching with 16 steps achieves slightly faster inference without loss of quality compared to diffusion policy with 16 steps. We hypothesize that flow matching with optimal transport generates straighter flows, and thus causes faster inference.
Fig. 1 showcases more examples of the waypoint motion path and vector field plots of flow matching policy. We could see that flow matching transforms random waypoints to the target trajectory from timestep 0 to 1. The animations in the supplementary video better demonstrate the whole process.
4.2.3 Comparisons against diffusion policy
@@
Fig. 7 shows the training and testing loss of flow matching and diffusion policy throughout the training process. We can see flow matching exhibits greater stability on both training and evaluation than diffusion model. For the Push-T benchmark evaluation, flow matching also requires fewer steps and better performance (120 steps, score 1.0) to push a T-shaped block to fully overlap with a fixed target using a circular end-effector, compared to diffusion policy (200 steps, score 0.91).
@@ 4.2.4 Ablations
Inference Steps: Table 2 also showcases better performances of flow matching when applying more inference steps, with a trade-off of longer inference time.
@@
Trajectory Horizon: We empirically test trajectory representation with 8, 16, 32 and 64 waypoints. More waypoints are not necessary, while fewer waypoints are unable to entirely encapsulate the complete long-horizon trajectories. We have found out 32-waypoint trajectories to be optimal. In general, the trajectory horizon does not wield a significant influence on flow matching performance.
4.3 Real-World Robot Experiments
Lastly, we deploy flow matching and other policies on real robot manipulation evaluation, as shown in Table 3. We carry out 50 replications of trials for each policy. The orientation is predefined. We use a KINOVA Gen3 arm, an Azure Kinect camera, and a 3D printed UMI gripper (Chi et al., 2024) for real-world robot experiments.
We observe that most failures occurred during the initial grasping. Even the trajectory prediction is accurate in general, a slight misalignment between the first trajectory waypoint and the object would cause grasping failures. This problem might be addressed in the future by a weighted flow matching policy, which highlights the initial waypoints by adding higher probability weights. Since we have not considered closed-loop actions, the robot might also collide with objects in the process of moving to the first waypoint. Some out-of-distribution factors also lead to failure, including background and camera view changes. Larger data sizes and more diversity should alleviate this problem.
We also investigate how the affordance would guide the flow matching policy. We trained flow matching for each task separately, taking the raw RGB images as input. It achieves a higher loss but a similar success rate compared to training all tasks in one policy with affordance guidance. Interestingly, a further comprehensive examination reinforces the argument that flow matching could handle multimodal action distribution. Fig. 8 shows an example. From the left figure, we can see that when training a separate policy of moving the towel toward the trash for sweeping, the predicted trajectory (yellow) could be detached from the ground truth (red), but still a reasonable solution that allows for a successful robot execution. With affordance guidance, the prediction is closely aligned with the truth (Fig. 8-right).
Figure 8: Examples of how visual affordance guides the flow matching policy.
Methods Flow Diffusion Transformer Flow Matching
Matching Policy BC (separate policy)
Success Rate 0.82 0.76 0.44 0.80
Table 3: Success rate of real-world robot experiments.

### v2

@@ 3.2.3 Intriguing Properties of Flow Matching Policy
     $\bm{x}_{t}=t\bm{x}_{1}+(1-t)\bm{x}_{0}$, linear interpolation
     $\mathbf{v}_{t}(\bm{x}|\bm{o})=v_{\bm{\theta}}(\bm{x}_{t},t|\bm{o})$, flow estimation
     $\nabla_{\bm{\theta}}\left\|v_{\bm{\theta}}(\bm{x}_{t},t|\bm{o})-\dot{\bm{x}}_{t}\right\|$, gradient step
@@
Stopping criteria: $1,000$ training epochs
Algorithm 2 Robot Diffusion Policy (DDPM) Training
observation $\bm{o}$, target robot actions $\bm{x}_{1}$, source Gaussian noises $p_{0}$
noise $\mathbf{\epsilon}_{\bm{\theta}}$
while not converged do
     $\bm{x}_{0}\sim p_{0}$, sample Gaussian noises
     $t\sim\mathcal{U}[0,1]$, sample time steps
     $\bm{x}_{t}=\mathcal{N}(\mathbf{x}_{t};\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0},(1-\bar{\alpha}_{t})\mathbf{I})$, forward process
     $\mathbf{\epsilon}_{t}(\bm{x}|\bm{o})=\epsilon_{\bm{\theta}}(\bm{x}_{t},t|\bm{o})$, noise estimation
     $\nabla_{\bm{\theta}}\left\|\epsilon_{\bm{\theta}}(\bm{x}_{t},t|\bm{o})-\bm{\epsilon}_{t}\right\|$, gradient step
end while
Stopping criteria: $1,000$ training epochs
3.3 Activities of Daily Living Dataset
We construct a real-world dataset with $10$ tasks across Activities of Daily Living to test the proposed method. Each task includes $1,000$ sets of RGB images, demonstrated robot trajectories, and labeled ground truth of affordances. Thus $10,000$ demonstrations have been collected in total. The data has been manually collected by moving robot end-effectors with kinesthetic teaching. We label the heatmap affordance of our data with 2D Gaussian blobs centered on the pixel of demonstrated action during data collection. The novelty of our dataset could be summarized as: (i) It contains the same scenarios with multiple objects, multi-task affordances, and the demonstrated robot trajectories. (ii) All tasks are related to Activities of Daily Living that involve (simulated) human data.
Objects are randomly placed within the range of the table (1.5 meters $\times$ 1 meter). The human (manikin) is placed randomly around the table in the camera view. We have around 30 different objects. The camera view is fixed as we have not used 3D representation, which will be addressed in our future work.
Our tasks include prompt primitives: ‘sweep the trash’, ‘pass the water to the human’, ‘hang the towel’, ‘put on the hat’, ‘cover the food’, ‘wipe the nose’, ‘wipe the forearm’, ‘feed the human’, ‘comb the hair’, and ‘brush the teeth’. Each task corresponds to one prompt primitive. We could also add a pretrained LLM layer (e. g., GPT) in the front for zero-shot text classification, allowing for linking other language instructions to one of the ten prompt primitives. The dataset will be made publicly available soon.
@@ 4 Experiments
We systematically evaluate prompt tuning method and flow matching policy against baseline studies. We also investigate how design choices would affect their performance. We randomly split the Activities of Daily Living dataset with $80$%-$20$% percentage of training and testing. The results reported here are obtained after $1,000$ epochs of training with 3-fold cross validation to minimize the impact of any noises.
@@ 4.1 Affordance Evaluation with Prompt Tuning
4.1.1 Baseline Studies
For a fair comparison, all the baselines here use self-supervised pretrained MAE weights on ImageNet-21k dataset for the vision transformer model. In ablation studies, we will investigate different pretrained weights. The training parameters for the prompt tuning network include an image size of 224 × 224, AdamW with a learning rate of 1.5e-4 including Warmup with step-decay, and batch size of 256. The text encoding layers output 76 tokens, as the CLIP method (Radford et al., 2021).
4.1.2 Main Results
@@
Table 1 presents the results of prompt tuning on our testing dataset for affordance learning, comparing against baselines. We use two metrics to evaluate our results: (i) L2 error of affordance heatmap estimation (the fourth column), and (ii) L2 distance between the predicted and ground truth of heatmap centers (the fifth column). For the inference, we fit Gaussian Mixture Models to determine the heatmap centers. The heatmap error is averaged on each map, and the center error is averaged on per center point.
Four observations could be made from this result:
  1. 1.
    The deep structure outperforms the other baselines except for the full finetuning.
  2. 2.
    Full finetuning slightly outperforms deep prompt tuning in terms of heatmap estimation error and heatmap center error. However, the distinction of heatmap center errors ($1.78$ pixels) remains subtle, given the full image size of $224\times 224$. This outcome is favorable as it indicates that most heatmap errors are caused by the tails of the Gaussian distribution, instead of the center area where the robot actions actually applied on.
  3. 3.
    As expected, the shallow structure performs suboptimally to the deep structure. In the following ablation studies, we will investigate how the prompt depth affects the performance.
  4. 4.
    Generalizability: We also observe that the trained model could be generalized to new objects. For example, the training dataset only includes a manikin. We found out that it generates well on our testing data with real humans. Note that the proposed tuning method is parameter-efficient, it is envisaged that the method could be readily transferred to different tasks with a small amount of task-specific data.
@@
What do prompts learn? We show a t-SNE (Van der Maaten and Hinton, 2008) visualization of the embeddings after the last vision transformer layer and before the decoder in Fig. 5. We can see that the points of the same color (e. g., tasks with the same language prompts) are embedded together, implying that the representations recover the underlying manifold structure of discriminative task information.
@@ 4.1.3 Ablations
We are aware of other variants of vision transformers or weights, for example, CLIP vision encoder. As our goal is to integrate textual representations into any vision encoder while keeping it frozen, preserving its visual understanding capabilities, we have chosen the most basic ViT-B-16 transformer backbone and commonly used pretrained weights. We have also tried GPT directly, but no constant performance was achieved without carefully designed prompts or finetuning.
@@ 4.2.1 Baseline Studies
We compare our flow matching policy against two other robot behavior cloning methods: (i) Diffusion Policy (Chi et al., 2023) with Denoising Diffusion Probabilistic Models (DDPM), and (ii) Transformer-based behavior cloning with Mean Square Error Loss, as customary in RVT (Goyal et al., 2023), RT-X (Padalkar et al., 2023).
Note that we are aware of other well-performed robot behavior cloning methods, including energy-based IBC (Florence et al., 2022), GAIL (Ho and Ermon, 2016), etc. Since extensive studies have been conducted and showed lower performance of these methods compared to diffusion policy, we choose two representative baselines for evaluation.
We consider CNN-based and Transformer-based structures for flow matching. The flow matching and baselines are trained in a supervised manner using RGB images with visual affordances as input, and the output is a long-horizon trajectory in both 2D pixel space and 3D Cartesian space. All 10 tasks in our dataset are trained in a single policy. Table 4 shows the hyperparameters we have used in flow matching and diffusion policy. Table 5 shows the tasks Summary.
@@
Baselines Diffusion Policy (2-step inference) 38.37 25.08 24.81
Diffusion Policy (16-step inference) 0.884 2.096 159.72
@@
Flow Matching (CNN, 1-step inference) 0.898 2.151 8.53
Ablations Flow Matching (CNN, 2-step inference) 0.888 2.104 24.29
Flow Matching (CNN, 4-step inference) 0.846 2.094 44.99
Flow Matching (CNN, 8-step inference) 0.842 2.092 70.10
Flow Matching (CNN, 16-step inference) $0.841$ $2.091$ 100.98
@@ 4.2.2 Main Results
Table 2 presents the results of flow matching policy on our testing dataset for robot trajectory learning, comparing against baselines. We use two metrics for evaluation: (i) the error of 2D and 3D trajectory estimation, and (ii) the average inference time across various steps, performed using PyTorch with RTX 4090 GPU acceleration. The trajectory error is averaged on each point of the trajectory.
Figure 7: Training and testing loss of flow matching and diffusion policy throughout the training process. Flow matching exhibits greater stability on both training and evaluation than diffusion policy.
Three observations could be made from this result:
  1. 1.
@@
    Generation Quality: Flow matching (CNN-based, 16 steps) outperforms other baselines in terms of 2D and 3D trajectory prediction accuracy. The Transformer behavior cloning achieves marginal precision. This is expected as it is hindered by the nature of multi-modal action distribution, causing the averaging out across non-convex spaces.
  2. 2.
@@
    Stability: Fig. 7 shows an example of training and testing loss of flow matching and diffusion policy throughout the training process. We can see flow matching exhibits greater stability on both training and evaluation than diffusion policy.
  3. 3.
    Inference Time: In terms of inference time, Table 2 showcases that flow matching with 16 steps achieves faster inference time compared to diffusion policy with 16 steps. We hypothesize that flow matching with linear pointwise flows (following the optimal transport theory) generates straighter flows, and thus causes faster inference. Table 2 also showcases better performances of diffusion policy when applying more inference iterations, with a trade-off of longer inference time. Contrarily, flow matching has not shown significant improvements when increasing inference steps. We hypothesize that this is because flow matching trains Continuous Normalizing Flow models, where an ordinary differential equation is solved without learning a series of discrete steps to progressively refine the generated sample. This considerably reduces the inference time for closed-loop robot manipulation, as 2-step flow matching (error: $0.888$cm, time: $8.53$ms) has achieved comparable performance as 16-step diffusion policy (error: $0.884$cm, time: $159.72$ms), but prominently lower inference time roughly by 85%.
4.2.3 Real-World Robot Experiments
We deploy flow matching and other policies on real robot manipulation evaluation, as shown in Table 3. We carry out 50 replications of trials for each policy. The orientation is predefined. We use a KINOVA Gen3 arm, an Azure Kinect camera, and a 3D printed UMI gripper (Chi et al., 2024) for real-world robot experiments.
We observe that most failures occurred during the initial grasping. Even the trajectory prediction is accurate in general, a slight misalignment between the first trajectory waypoint and the object would cause grasping failures. This problem might be addressed in the future by a weighted flow matching policy, which highlights the initial waypoints by adding higher probability weights. Some out-of-distribution factors also lead to failure, including background and camera view changes. Larger data sizes and more diversity should alleviate this problem.
RGB images ablations against affordances: We also investigate how the affordance would guide the flow matching policy. We trained flow matching for each task separately, taking the raw RGB images as input. It achieves a higher loss but a similar success rate compared to training all tasks in one policy with affordance guidance. Interestingly, a further comprehensive examination reinforces the argument that flow matching could handle multimodal action distribution. Fig. 8 shows an example. From the left figure, we can see that when training a separate policy of moving the towel toward the trash for sweeping, the predicted trajectory (yellow) could be detached from the ground truth (red), but still a reasonable solution that allows for a successful robot execution. With affordance guidance, the prediction is closely aligned with the truth (Fig. 8-right).
Figure 8: Ablations between using RGB images and affordances as inputs. We show examples of how visual affordance guides the flow matching policy.
@@ 4.2.4 Ablations
Trajectory Horizon: We empirically test trajectory representation with 8, 16, 32 and 64 waypoints. More waypoints are not necessary, while fewer waypoints are unable to entirely encapsulate the complete long-horizon trajectories. We have found that, in general, the trajectory horizon does not wield a significant influence on flow matching performance. The performance reported in the above main results section is achieved by using 32 waypoints to represent the trajectory.
4.3 Comparisons between flow matching policy diffusion policy
To further investigate the performance of flow matching compared to diffusion policy, we benchmark the proposed methods on two more datasets: (i) Push-T adapted from (Florence et al., 2022); and (ii) Franka Kitchen proposed in Relay Policy Learning (Gupta et al., 2019).
  1. 1.
    Push-T requires pushing a T-shaped block to a fixed target with a circular end-effector. Push-T takes RGB images with proprioception of end-effector location as inputs, and end-effector actions as outputs. The experiments are carried out in a closed-loop manner. All methods are trained with $200$ demonstrations for $3,000$ epochs. Variation is added by random initial conditions for T block and end-effector.
  2. 2.
    Franka Kitchen contains 7 objects for interaction and comes with a human demonstration dataset of $566$ demonstrations, each completing 4 tasks in arbitrary order. The goal is to execute as many demonstrated tasks as possible, regardless of order, showcasing both short-horizon and long-horizon multimodality. The training takes states as inputs, and robot joint actions as outputs for $4,500$ epochs. This benchmark includes 6D closed-loop robot position and orientation actions, as well as gripper actions. Variation is added by random initial conditions for the robot and object states.
    For Push-T and Franka Kitchen benchmarks, we report results from the average of the last checkpoint with $10$ trials of replication. In addition, we also report the best perforce in the replications of the last checkpoint. The evaluation is carried out across $500$ different environment initial conditions ($5,000$ in total).
  3. 3.
    Real-world Daily Living Task, for clarity, takes affordance images as inputs, and 3D position end-effector trajectories as outputs. One-shot open-loop estimation of the entire trajectory is used in our dataset. We randomly split the dataset with $80$%-$20$% percentage of training and testing. The results reported are obtained after $1,000$ epochs of training with 3-fold cross validation to minimize the impact of any noises. Other training and testing details have been illustrated in the above sections.
Table 4 shows the hyperparameters we have used in flow matching and diffusion policy. Table 5 shows the tasks Summary.
Similar conclusions could be achieved from Table 3, as in the Main Results section: (i) Generation Quality: Flow matching outperforms diffusion policy in all benchmarks of robot manipulation experiments. (ii) Inference Time: Flow matching considerably reduces the inference time for closed-loop robot manipulation. For example, in Franka Kitchen benchmark, 2-step flow matching (success rate: $0.9750$, time: $13.18$ms) has achieved comparable performance as 16-step diffusion policy (success rate: $0.9840$, time: $157.46$ms), but prominently lower inference time roughly by 90%.
Methods Flow Matching Diffusion Policy Flow Matching Diffusion Policy Transformer Flow Matching
(Inference Step) ($2$-step) $\uparrow$ ($2$-step) $\uparrow$ ($16$-step) $\uparrow$ ($16$-step) $\uparrow$ BC $\uparrow$ separate policy
Push-T 0.8771/0.7111 0.4412/0.1872 $0.9035/0.7490$ 0.8840/0.7178
Franka Kitchen 0.9750/0.6134 0.2355/0.0527 $0.9960/0.7172$ 0.9840/0.6716
Real-world Daily Living Task $0.82$ 0.76 0.44 0.80
Table 3: We present the robot performance with different checkpoint selection methods in the format of (max performance) / (average of last checkpoint with $10$ trials of replication), with each averaged across $500$ different environment initial conditions ($5000$ in total) for testing. The metric used here is success rate, except for the Push-T task which uses target area coverage.
DDPM Ta Tp ImgRes D-Net D-Para V-Enc V-Para Lr WDecay Iters Train Iters Eval
Push-T 8 16 96x96 ConditionalUnet1D 73 ResNet-18 12 1e-4 1e-6 16 2/4/8/16
Franka Kitchen 8 16 N/A ConditionalUnet1D 73 N/A N/A 1e-4 1e-6 16 2/4/8/16
Real-world Daily Living Task 32 32 224x224 ConditionalUnet1D 73 ResNet-18 12 1e-4 1e-6 16 2/4/8/16
Flow Matching Ta Tp ImgRes F-Net F-Para V-Enc V-Para Lr WDecay Iters Train Iters Eval
Push-T 8 16 96x96 ConditionalUnet1D 73 ResNet-18 12 1e-4 1e-6 N/A 2/4/8/16
Franka Kitchen 8 16 N/A ConditionalUnet1D 73 N/A N/A 1e-4 1e-6 N/A 2/4/8/16
Real-world Daily Living Task 32 32 224x224 ConditionalUnet1D 73 ResNet-18 12 1e-4 1e-6 N/A 2/4/8/16
Table 4: Hyperparameters for flow matching and Diffusion Policy. Ta: action horizon. Tp: action prediction horizon. ImgRes: environment observation resolution. D-Net: diffusion network. D-Para: diffusion network number of parameters in millions. F-Net: flow matching network. F-Para: flow matching network number of parameters in millions. V-Enc: vision encoder. V-Para: vision encoder number of parameters in millions. Lr: learning rate. WDecay: weight decay. Iters Train: number of training diffusion iterations. Iters Eval: number of inference iterations. Network structures for flow matching policy are modified in the same way in diffusion policy.
Tasks Rob Obj ActD PH Steps Img Closed-loop
Push-T 1 1 2 200 300
Franka Kitchen 1 7 9 566 280
Real-world Daily Living Task 1 $\approx 30$ 3 8,000 N/A
Table 5: Tasks Summary. Rob: number of robots. Obj: number of objects. ActD: action dimension. PH: proficient-human demonstration. Steps: max number of rollout steps.